OpenStudy (anonymous):

please help me factor: x^2+2xy+2x+y^2+2y-8

OpenStudy (tkhunny):

Why do you believe you can factor it?

OpenStudy (anonymous):

it's on a worksheet

OpenStudy (tkhunny):

Fair enough. What's your plan?

OpenStudy (anonymous):

i have the answer key, but when i do it. i get it wrong

OpenStudy (anonymous):

x^2+2xy_y^2 factors to (x+y)^2 and then factor 2 out of 2x+2y-8...but it;s after that

OpenStudy (anonymous):

it's wrong*

OpenStudy (tkhunny):

Okay, then why not do it in a way that works? As in, try something else.

OpenStudy (anonymous):

on the answer key, it only factored out 2 for 2x+2y only? i dont know why. Can you explain to me?

OpenStudy (tkhunny):

If we are going to get x^2, y^2, xy and a constant term, it has to factor like this: (ax+by+c)(dx+ey+f) Since we'll need a coefficient of 1 on x^2, then a = d = 1 (x+by+c)(x+ey+f) Since we'll need a coefficient of 1 on y^2, then b = e = 1 (x+y+c)(x+y+f) By a little luck, maybe, this also produces 2xy. Since we'll need a constant term of -8, then c*f = -8 The coefficients on x and y suggest also that c + f = 2 Can you solve for c and f?

OpenStudy (tkhunny):

If you like, we can treat it like any other quadratic expressions. x^2 + 2xy + y^2 + 2x + 2y - 8 (x^2 + 2xy + y^2) + (2x + 2y) - 8 (x + y)^2 + 2(x + y) - 8 If this said, z^2 + 2z - 8, could you factor it? ((x+y)+4)((x+y)-2)